You have 63 minutes to take an exam of 24 questions. After you finish the first 11 questions, you realize you must answer the remaining questions twice as fast to finish the exam. How many questions per minute were you answering? Answer in units of questions/ minute.
Answer by lwsshak3(11628)

You have 63 minutes to take an exam of 24 questions. After you finish the first 11 questions, you realize you must answer the remaining questions twice as fast to finish the exam. How many questions per minute were you answering? Answer in units of questions/ minute.
***
let x=minutes took to finish first 11 questions
11/x=rate(questions/min )took to finish first 11 questions
22/x=rate at which you must finish remaining 13 questions
63-x=remaining minutes to finish 13 questions
..
22/x=13/(63-x)
13x=1386-22x
35x=1386
x=39.6 min
rate=11/39.6=0.28 (questions/min)
How many questions per minute were you answering? 0.28